40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Tue Sep 3 16:48:23 2019, anchal, DailyProgress, NoiseBudget, Daily updates on noise budget 
    Reply  Wed Sep 11 10:42:50 2019, anchal, DailyProgress, NoiseBudget, 500 Hz hump gone! 
       Reply  Mon Sep 30 16:47:15 2019, anchal, DailyProgress, NoiseBudget, Adding links again; Comparison with Old Estimates on same plot 
Message ID: 2424     Entry time: Wed Sep 11 10:42:50 2019     In reply to: 2409     Reply to this: 2437
Author: anchal 
Type: DailyProgress 
Category: NoiseBudget 
Subject: 500 Hz hump gone! 

blush

Checkout the latest noisebudget. 500 Hz peak is gone now. This could be because of any of the changes I made in past 2 days:

1) NPMC servo gain increased. I changed R8 from 9.09k to 51 Ohms. That increased the gain by a factor of 200 which is equivalent to 46 dB. The slider for AD602 gain decreased by 23 dB to remain below oscillations wiht a 1dB gain margin. So maybe this reduced the noise a lot more (even though it is still present)

2) I did replace a normal beam dump (91,45) on the reflection path from North PMC with a triangular black glass cavity and I oriented another one (94,30) just before NPMC blocking rejection from a PBS as it was oriented horribly wrong.

3) We also changed the detector to NF1811, but that hasn't resulted in changing this feature in the past.

CTN Daily Beatnote Spectrum

CTN Latest Beatnote Spectrum


Note that the expected noise has now decreased roughly by 3 times in the range 200 Hz to 500 Hz. See CTN: 2423

ELOG V3.1.3-